OpenStudy (anonymous):

evaluate the integral log 10 x/x dx

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\log_{10}x}{x} dx ?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

you remember change of base formula \[\log_{10}x=\frac{lnx}{\ln10}\] \[\int\limits_{}^{}\frac{lnx}{\ln10}*\frac{1}{x} dx\] let u=lnx du=1/x dx \[\frac{1}{\ln10}\int\limits_{}^{}u du\]

myininaya (myininaya):

\[\frac{1}{\ln10}*\frac{u^2}{2}+C=\frac{1}{2\ln10}(lnx)^2+C\]

OpenStudy (anonymous):

i did not understand the last part?

myininaya (myininaya):

the substitution?